Calculate the ∆G for the dissociation of acetic acid in a 5.76 pH buffer at 298 K. Now what do you expect acetic acid to do in this higher pH buffer. What kind of answer do you expect G to be? Positive or negative?
Answer #1

as per the relationship eqation between

delta G and reaction equilibrium constant Keq

∆G = - RT ln Keq. -----------(1)

G is free energy , R is gas constant , T is obsolute temperature now we know that Keq = [products concentration]/ [Reactants]

Keq = [ buffer concentration ] = 10 ^ ( - pH ) = 10^ ( - 5.67 )

Keq = 2.14 x 10 ^-6

now putting all values in eqation 1

∆G = - 8.314 x 298 x ln ( 2.14 x 10^-6) =

= - 2477.6 x ln (2.14 x10^-6) = - 2477.6 x ( -13.054)

= 32344.34 joules

∆G = 32.34434 kilo joules

delta G is POSITIVE